# The ubiquity of variable radio emission and spin-down rates in pulsars

### v1

@@ 4 Results and analysis
    long-term smooth variations in $\dot{\nu}$ over long timescales,
@@
    and transient increases or decreases in $\dot{\nu}$.
Profile variability maps were generated for 242 pulsars, which can be found in the Supplementary Materials. An example of one is shown in Figure 7 for PSR J1048--5832, where the upper and lower panels present the varying spin spin-down rate and profile shape changes respectively. The remaining 17 of the 259 pulsars in our full sample had insufficient per-epoch $S/N$ to generate accurate variability maps. From these maps, we found that 52 pulsars showed substantial changes in profile shape. We also identified several pulsars that appear to show intense epoch-to-epoch shape variations across the entire profile that are approximately symmetric about the peak. This behaviour can be attributed to profile jitter, resulting from the finite number of rotations that are recorded in our relatively short, per-pulsar observations. Pulsars that were most strongly affected by this include PSRs J0837--4135, J1430-6623 and J1644-4559, which are not included in the profile shape change analysis.
Figure 4: Spin-down timeseries for the 238 pulsars that displayed significant variability. Labels on the left-hand side of each panel indicates the pulsar J2000 name and the percentage difference between the minimum and maximum value of $\dot{\nu}$.
@@ 4.1 Known variable pulsars
PSR J0742--2248 (B0740--28)
Among the initial sample of variable profile/spin-down pulsars reported by Lyne et al. (2010), PSR J0742--2248 has a fast variability timescale of only $\sim 135$ d. The relationship between changes in profile shape and spin-down rate of this pulsar is complicated. Both are largely uncorrelated on long-time scales, but display an increased levels of correlation following the MJD 55022 glitch (see, Keith et al. 2013), suggested to be a result of glitch-induced coupling between the internal and external dynamics of the pulsar. This behaviour is readily apparent in the variability map as deficits (excess) emission from the leading (dip between leading and trailing) profile components. The increased correlation appears to vanish $\sim$330 d before the MJD 56727 glitch, and fails to re-appear in the post-glitch spin-down/shape variations. Similarly, the MJD 59839 glitch also does not result in enhanced correlation between the profile and spin-down rate, indicating the phenomenon identified by Keith et al. (2013) is not a universal feature of the post-glitch behaviour of PSR J0742--2248.
@@
A 0.33 s pulsar that has been reported to switch between two emission states: ‘mode A’ where the emission occurs slightly earlier in pulse phase, and ‘mode B’ where it is both weaker and arrives slightly later (Zhang et al., 2020). Our spin-down model displays rapid, short-duration decreases in spin-down rate with a quasi-periodic spacing of $\sim$268 d. These spin-down events typically last between 130-200 d with typical amplitudes of $\delta\dot{\nu}/\dot{\nu}\sim 1.58$%. The profile variability map shows the trailing edge of the profile displays changes on an equally rapid timescale, usually appearing as an excess of emission whenever the pulsar is in the lower spin-down state that is only mildly correlated with the spin-down timeseries. An excess of emission in this part of the profile would point to the pulsar preferentially emitting in ‘mode B’ between the quasi-periodic decreases in spin-down rate.
@@
Rocketing away from the putative supernova remnant G270.4--1.0 (Johnston & Lower, 2021), this orthogonally rotating pulsar was found by Brook et al. (2016) to display correlated changes in the leading edge of the main pulse and both inter-pulse components. All three of these profile components display simultaneous excesses and deficits of emission, which are also reflected in our updated profile variability map. Our profile/spin-down correlation map shows a positive correlation between the changes in these profile components that lag corresponding spin-down variations by $\sim$300 d, while the more rapid changes in the peak of the main-pulse being uncorrelated with spin down rate. Similar cross-talk between emission originating from antipodal magnetic poles was previously identified in the single-pulse mode switching of PSRs J1057--5226, J1705--1906 and J1825--0935555Note that model fits to the linear polarisation position angle swing of PSR J1825--0935 are consistent with both an aligned or orthogonal rotator (Johnston et al., 2023). (Weltevrede et al., 2007, 2012; Fowler & Wright, 1982), indicating emission state-switching is a global phenomenon that affects the plasma content of the entire magnetosphere. We also detect more rapid variations in the primary peak of the main-pulse, though this may be the result of longitudinal profile jitter. A small glitch occurred in PSR J0908--4913 on MJD 58765 (Lower et al., 2021), which does not appear to have had a noticeable impact on the post-glitch profile shape or spin-down variability.
@@
Figure 7: Rotation and emission changes in the known variable pulsar PSR J1048--5832. Top panel shows the spin-down timeseries, lower panel the profile variability map with the median profile depicted on the right. Orange dashed lines in the top panel indicate the epochs of glitch events.
A bright gamma-ray pulsar, PSR J1048--5832 is known to display quasi-periodic switching between a weak and strong emission state every $\sim$17 rotations (Yan et al., 2020). Our spin-down model and profile variability map in Figure 7 shows this variable emission behaviour also extends to longer timescales. The profile variations appear confined to the leading shoulder and primary profile components, each displaying continuous evolution with time similar what is observed in the main peaks of PSRs J0908--4913, J1600--5400 and J1602--5100. It is therefore not surprising that the time-averaged inter-glitch profiles highlighted by (Liu et al., 2024) appear different to one another. The spin-down variations are qualitatively complex, though a Lomb-Scargle periodogram recovers a strong quasi-period at $\sim$480 d. Our observed changes in the profile shape are only weakly correlated with $\dot{\nu}$ on a similar timescale.
@@
An orthogonal rotator, PSR J1057--5226 is known to display simultaneous, short-timescale emission state switching across both its main pulse and inter pulse (Biggs, 1990; Weltevrede et al., 2012). The spin-down rate of this pulsar varies smoothly over time by $\delta\dot{\nu}/|\dot{\nu}|\sim 0.27\%$, with no clear quasi-periodicity picked up in the Lomb-Scargle periodogram. Our profile variability map reveals inter-observation changes in the pulse profile of this pulsar are largely restricted to the main pulse, where the brightest peak displays rapid fluctuations in intensity, while the more central component has a long-term decay in intensity superimposed on similar rapid fluctuations. This slow variation is strongly anti-correlated with the spin-down timeseries.
@@
Timing of this pulsar at HartRAO over 13 yr revealed its timing residuals displayed oscillatory variations with a best-fit quasi-period of 2560 d (Frescura & Flanagan, 2003). Gaussian process regression analysis of the HartRAO timing by Maritz et al. (2015) showed this behaviour can be explained by quasi-periodic spin-down variations occurring on this timescale. Our analysis of the Murriyang observations clearly recovers the long-term quasi-periodic changes in spin-down rate of PSR J1326--5859. However, the periodogram reveals a sharp quasi-period of only 1272 d, approximately half the period found by Frescura & Flanagan (2003). This pulsar also displays significant profile shape changes, though the high amplitude, rapid variations either side of the main peak are likely a result of profile jitter as an excess on one side of the peak is always accompanied by a deficit on the other. The more subtle long-term changes in the precursor and tail components appear to be genuine state changes in the pulsar emission. The precursor component changes are weakly correlated with variations in the spin-down rate at zero-lag while changes in the tail lag the spin-down rate by $\sim 500$ d.
@@
Identified by Brook et al. (2016) as a variable pulsar, the Gaussian-like profile of PSR J1359--6038 often displays periods of excess emission on the right-hand slope that are correlated with transient decreases in spin-down rate that gradually decay back toward a steady state. An extreme broadening of both the left and right hand sides of the profile coinciding with a sudden $\sim 0.14\%$ drop in spin-down occurred between MJD 56450-56650 (Brook et al., 2016). Despite the much larger overall alteration in pulse shape, the spin-down rate change associated with this event is of comparable amplitude to those that coincide with the less substantial profile shape changes. Our periodicity analysis of the spin-down timeseries reveals a strong quasi-period at 361 d, which is likely an artefact of a small offset in the pulsar position.
@@
This is a $0.3$ s orthogonal rotator that was found to display ‘on/off’ sub-pulse modulation across both its the main and inter pulse (Weltevrede et al., 2007). Like PSR J1057--5226, the spin-down rate of PSR J1705--1906 varies slowly over time, with an apparent, dominant quasi-period of $\sim$2645 d. Our profile variability map does not recovery any shape variations that are substantially different from fluctuations in the off-pulse noise, suggesting the rapid state switching every 10.4 rotations identified by Weltevrede et al. (2007) must be stable over long time scales. This is unlike other state switching pulsars, where the profile shape changes recovered in the profile variability maps reflect the pulsar preferentially spending more time in one emission state than the other.
@@
This $1.4$ s pulsar was noted by Johnston et al. (2021) as switching between two distinct emission states: a ‘weak’ state where the profile consists of three approximately equal intensity sub-components and a ‘bright’ mode where the central component is up to an order of magnitude brighter than in the weak-state. Our spin-down model displays small variations with time of up to $\delta\dot{\nu}\sim 0.14\%$, albeit with large uncertainties, that is only weakly correlated with changes in profile shape. A small glitch with an amplitude of $\Delta\nu_{g}/\nu=11.21(5)\times 10^{-9}$ and change in spin-down rate of $\Delta\dot{\nu}_{g}/\dot{\nu}=3.0(2)\times 10^{-3}$ occurred around MJD 58570. Curiously, the profile variability map shows the bright state appears to have been more prevalent in observations taken before MJD 58750. It is unclear whether the two phenomena are related since the bright-mode deficit lags the glitch by $\sim 100$ d.
@@
Identified by Basu et al. (2024) as displaying alternating fluctuations in intensity between the two peaks of its profile in long-term monitoring observations with MeerKAT. This behaviour appears to be a variant of that identified in the Lyne et al. (2010) sample, where the apparent switching between profile component intensities is due to the pulsar preferentially emitting from one component over the other (see the single-pulse data in Figure 5 of Basu et al. 2024). The switching in component intensities is marginally visible in our profile variability map. Our spin-down timeseries reveals the pulsar initially displayed clear quasi-periodic variability with two peaks in the Lomb-Scargle periodogram at 157 d and 314 d. This behaviour was seemingly interrupted by a glitch that occurred around MJD 58236, after which the quasi-periodicity becomes less defined, which may partially explain the lack of correlation with the emission state changes.
@@
The main peak and trailing shoulder components of this pulsar were noted by Basu et al. (2024) as displaying correlated fluctuations in intensity, where an increases in the emission peak corresponded to a deficit in trailing shoulder emission, and vice-versa. No spin-down variability was reported in their analysis of PSR J1741--3927, however this appears to have been a result of the short timing baseline of the MeerKAT Thousand Pulsar Array, as our $\dot{\nu}$ timeseries shows a clear long-term fluctuation over our 10 yr of Murriyang observations. The profile variability map that we generated also shows the main peak undergoing clear, rapid variations on a timescale of a few hundred days, though we do not see similar anti-correlated changes in the trailing shoulder component. Instead, both the leading and trailing shoulders appear to undergo a secular increase in intensity that is correlated with the long-term change in $\dot{\nu}$ at a 1000 d lag.
@@ 4.2 New pulsars with both profile and spin-down variations
Alongside the known variable pulsars, we have discovered another 28 that show fluctuations in both profile shape and spin-down rate. Here we summarise the kind of variations detected in the radio emission and spin-down properties of these pulsars.
@@
This pulsar displays a broadly Gaussian shaped radio profile, albeit with many blended sub-components. Its spin-down timeseries in Figure 4 exhibits slow $\delta\dot{\nu}/\dot{\nu}\sim 0.72$% fluctuations over 23.5 yr of timing. This is contrasted by the profile variability map, which shows rapid changes across much of the profile peak that appear somewhat uniform across the smaller sub-components. Our cross-correlation analysis returns no substantial positive or negative correlation between the profile variability map and the spin-down timeseries.
@@
Similar to PSR J0742--2248, this pulsar displays a complicated relationship between its highly variable profile shape and spin-down rate. The spin-down rate shows rapid changes with time, with quasi-periods of 456 d and 177 d. The entire profile is affected by shape variations that appear strongest across the peak, which also vary on short timescales. Emission and spin-down processes appear to be uncorrelated with one another similar to the pre-glitch behaviour of PSR J0742--2248.
@@
Another pulsar that displays rapid, albeit seemingly uncorrelated, variations in both its radio profile shape and spin-down rate over time. The profile shape changes appear in both components of the blended, two-peaked profile with no clear quasi-stable states that the pulsar switches between. Given the rapid changes in both profile shape and spin-down, it is possible our approximately monthly observing cadence is unable to fully resolve any underlying quasi-stable states. These somewhat random profile shape changes are interspersed with brief periods of enhanced emission from a bump in the trailing shoulder, the two most prominent of which occurred around MJD 58915 and MJD 59768.
@@
This pulsar displays slow changes in spin-down rate with an amplitude of $\delta\dot{\nu}/\dot{\nu}\sim 0.12$%. However its variable profile shape fluctuates on a much shorter month-long timescales. The observed profile variations appear as increases in width and are most prominent in the trailing edge of the overall Gaussian shape. These changes in the trailing edge are moderately correlated with the spin-down timeseries. Inspection of the profile variability map shows the enhanced emission from this part of the profile appears more prevalent with the pulsar is in a lower spin-down state.
@@
This is a 0.18 s pulsar that has undergone four glitches since the beginning of our timing programme in 2007. The spin-down timeseries of PSR J1453--6413 is largely flat, except for two noteworthy fluctuations in $\dot{\nu}$, one that took place between MJD 57057-58645 and another that began around MJD 59182 and is ongoing. These events appear to be associated with excess emission from the leading bump in the pulse profile, though our correlation analysis suggests there is only a weak positive correlation between the variability map at these pulse phases and the spin-down timeseries. The more rapid profile shape changes that occur in the leading edge of the profile peak are uncorrelated with the changes in $\dot{\nu}$. It is unclear whether these profile/spin-down events are related to the small amplitude glitches that took place several hundred days prior to their onset.
@@
An example of a pulsar where changes in profile shape are not associated with variations in $\dot{\nu}$. PSR J1456--6843 has a complicated, albeit symmetric pulse profile that appears comprised of multiple blended components. Our variability map shows these components each varying randomly over time, which according to the correlation map is decoupled from the substantially slower $\delta\dot{\nu}/\dot{\nu}\sim 0.75$% change in spin-down rate over time.
@@
A triple component pulsar, with low intensity precursor and trailing components and a bright central component that has a ‘shoulder‘ on its trailing edge. All three components display shape variations different timescales. The precursor and trailing components both show simultaneous excess and suppressed emission that is correlated with the changing spin-down rate of the pulsar. Emission from the leading edge of the central component varies on a similar timescale, but lags changes in the precursor/trailing components by $\sim$400 d. The shoulder on the central component displays more rapid variations, yet appears to preferentially emit in a weak state in concert with the precursor and trailing components. Changes across the central component peak are anti-correlated with the varying spin-down rate and are out of phase with changes in the other components by half a cycle. While the spin-down rate of PSR J1559--4438 visually appears to display both long and short timescale variations, with the more rapid changes appearing quasi-periodic, no substantial peaks are detected in the corresponding Lomb-Scargle periodogram.
@@
A bright gamma-ray pulsar with a 0.37 s spin period, PSR J1745--3040 has a two-component radio profile with the brighter trailing component consisting of a bright peak and a leading bump. Both components display substantial epoch-to-epoch changes in intensity in the profile variability map that are uncorrelated with the slow, secular changes in $\dot{\nu}$ over time. This variability appears to be a result of both our short observation times and emission state switching in the single pulses of PSR J1745--3040, behaviour that is also seen in similarly bright pulsars such as PSRs J1048--5832 and J1243--6423. In Figure 10 we present a stack of single pulses detected by Murriyang, where the pulsar clearly switches between on and off emission states. When the pulsar is in the ‘on‘ state, we detect radio emission from both profile components, occasionally interspersed with very bright bursts from the main peak of the second component. Whereas the ‘off‘ state appears to be extended nulls, where no radio emission is detected down to the sensitivity limit of the instrument. A simple comparison between the mean on-pulse flux density and the off-pulse RMS reveals an approximate nulling fraction of 0.51 during the observation presented in Figure 10. Unlike the other pulsars among our sample that display similar behaviour, there does not appear to be a strong association between for the pulsar preferentially emitting in one emission state over the other and the changes in $\dot{\nu}$ over time. This may however, simply be due to the much lower frequency fluctuation timescale of $\dot{\nu}$ and relatively high nulling fraction of this pulsar.
@@
A pulsar with a blended two-component profile that displays rapid fluctuations in intensity confined to the brighter, leading peak. These intensity changes are mildly correlated with the slowly varying spin-down rate. There is a persistent emission deficit that coincided with the pulsar being in a high spin-down state between MJD 58122-59248.
@@
Similar to PSRs J0742--2248, this pulsar also displays both rapid variations in profile shape and spin-down rate with a low Q-factor quasi-period of $\sim$157 d. The shape changes affect the entire profile to some degree, but are strongest at the peak. Some of this variability may therefore be the result of profile jitter. However, the shape changes are correlated with the rapidly varying spin-down rate, albeit with a $\sim$50 d lag. No glitches have been detected in the timing of PSR J1752--2806 to date.

### v2

@@ 4 Results and analysis
    long-term smooth variations in $\dot{\nu}$ over long timescales and
@@
    transient increases or decreases in $\dot{\nu}$.
Profile variability maps were generated for 214 of pulsars that displayed significant $\dot{\nu}$ variations, which can be found in the Supplementary Materials. An example of one is shown in Figure 8 for PSR J1048--5832, where the upper and lower panels present the varying spin spin-down rate and profile shape changes respectively. The remaining 24 of the 238 variable pulsars had insufficient per-epoch $S/N$ to generate accurate variability maps. From these maps, we found that 52 pulsars showed substantial changes in profile shape. We also identified several pulsars that appear to show intense epoch-to-epoch shape variations across the entire profile that are approximately symmetric about the peak. This behaviour can be attributed to profile jitter, resulting from the finite number of rotations that are recorded in our relatively short, per-pulsar observations. Pulsars that were most strongly affected by this include PSRs J0837--4135, J1430-6623 and J1644-4559, which are not included in the profile shape change analysis.
Figure 5: Spin-down timeseries for the 238 pulsars that displayed significant variability. Labels on the left-hand side of each panel indicates the pulsar J2000 name and the percentage difference between the minimum and maximum value of $\dot{\nu}$. The downward pointing arrows indicate the epochs of detected glitches.
@@ 4.1 Known variable pulsars
PSR J0742--2822 (B0740--28)
Among the initial sample of variable profile/spin-down pulsars reported by Lyne et al. (2010), PSR J0742--2822 has a fast variability timescale of only $\sim 135$ d. The relationship between changes in profile shape and spin-down rate of this pulsar is complicated. Both are largely uncorrelated on long-time scales, but display an increased levels of correlation following the MJD 55022 glitch (see, Keith et al. 2013), suggested to be a result of glitch-induced coupling between the internal and external dynamics of the pulsar. This behaviour is readily apparent in the variability map as deficits (excess) emission from the leading (dip between leading and trailing) profile components. The increased correlation appears to vanish $\sim$330 d before the MJD 56727 glitch, and fails to re-appear in the post-glitch spin-down/shape variations. Similarly, the MJD 59839 glitch also does not result in enhanced correlation between the profile and spin-down rate, indicating the phenomenon identified by Keith et al. (2013) is not a universal feature of the post-glitch behaviour of PSR J0742--2822.
@@
This 0.33 s pulsar that has been reported to switch between two emission states: ‘mode A’ where the emission occurs slightly earlier in pulse phase, and ‘mode B’ where it is both weaker and arrives slightly later (Zhang et al., 2020). Our spin-down model displays rapid, short-duration decreases in spin-down rate with a quasi-periodic spacing of $\sim$268 d. These spin-down events typically last between 130-200 d with typical amplitudes of $\delta\dot{\nu}/|\dot{\nu}|\sim 1.58$%. The profile variability map shows the trailing edge of the profile displays changes on an similarly rapid timescale, usually appearing as an excess of emission whenever the pulsar is in the lower spin-down state that is only mildly correlated with the spin-down timeseries. An excess of emission in this part of the profile would point to the pulsar preferentially emitting in ‘mode B’ between the quasi-periodic decreases in spin-down rate.
@@
Rocketing away from the putative supernova remnant G270.4--1.0 (Johnston & Lower, 2021), this orthogonally rotating pulsar was found by Brook et al. (2016) to display correlated changes in the leading edge of the main pulse and both inter-pulse components. All three of these profile components display simultaneous excesses and deficits of emission, which are also reflected in our updated profile variability map. Our profile/spin-down correlation map shows a positive correlation between the changes in these profile components that lag corresponding spin-down variations by $\sim$300 d, while the more rapid changes in the peak of the main-pulse being uncorrelated with spin down rate. Similar cross-talk between emission originating from antipodal magnetic poles was previously identified in the single-pulse mode switching of PSRs J1057--5226, J1705--1906 and J1825--0935666Note that model fits to the linear polarisation position angle swing of PSR J1825--0935 are consistent with both an aligned or orthogonal rotator (Johnston et al., 2023). (Weltevrede et al., 2007, 2012; Fowler & Wright, 1982), indicating emission state-switching is a global phenomenon that affects the plasma content of the entire magnetosphere. We also detect more rapid variations in the primary peak of the main-pulse, though this may be the result of profile jitter shifting the exact peak of the profile from epoch-to-epoch. A small glitch occurred in PSR J0908--4913 on MJD 58765 (Lower et al., 2021), which does not appear to have had a noticeable impact on the post-glitch profile shape or spin-down variability.
@@
Figure 8: Rotation and emission changes in the known variable pulsar PSR J1048--5832. Top panel shows the spin-down timeseries, lower panel the profile variability map with the median profile depicted on the right. Orange dashed lines in the top panel indicate the epochs of glitch events. The short orange tick on the upper part of the variability map indicates the epochs of detected glitches, while the grey ticks along the bottom are the observation epochs at which the individual profiles were obtained.
This bright gamma-ray pulsar, PSR J1048--5832 is known to display quasi-periodic switching between a weak and strong emission state every $\sim$17 rotations (Yan et al., 2020). Our spin-down model and profile variability map in Figure 8 shows this variable emission behaviour also extends to longer timescales. The profile variations appear confined to the leading shoulder and primary profile components, each displaying continuous evolution with time similar what is observed in the main peaks of PSRs J0908--4913, J1600--5400 and J1602--5100. It is therefore not surprising that the time-averaged inter-glitch profiles highlighted by (Liu et al., 2024) appear different to one another. The spin-down variations are qualitatively complex, though a Lomb-Scargle periodogram recovers a strong quasi-period at $\sim$480 d. Our observed changes in the profile shape are only weakly correlated with $\dot{\nu}$ on a similar timescale.
@@
An orthogonally rotating pulsar, PSR J1057--5226 is known to display simultaneous, short-timescale emission state switching across both its main pulse and inter pulse (Biggs, 1990; Weltevrede et al., 2012). The spin-down rate of this pulsar varies smoothly over time by $\delta\dot{\nu}/|\dot{\nu}|\sim 0.27\%$, with no clear quasi-periodicity picked up in the Lomb-Scargle periodogram. Our profile variability map reveals that inter-observation changes in the pulse profile of this pulsar are largely restricted to the main pulse, where the brightest peak displays rapid fluctuations in intensity, while the more central component has a long-term decay in intensity superimposed on similar rapid fluctuations. This slow variation is strongly anti-correlated with the spin-down timeseries.
@@
Timing of this pulsar at HartRAO over a 13 yr revealed its timing residuals displayed oscillatory variations with a best-fit quasi-period of 2560 d (Frescura & Flanagan, 2003). Gaussian process regression analysis of the HartRAO timing by Maritz et al. (2015) showed this behaviour can be explained by quasi-periodic spin-down variations occurring on this timescale. Our analysis of the Murriyang observations clearly recovers the long-term quasi-periodic changes in spin-down rate of PSR J1326--5859. However, the periodogram reveals a sharp quasi-period of only 1272 d, approximately half the period found by Frescura & Flanagan (2003). This pulsar also displays significant profile shape changes, though the high amplitude, rapid variations either side of the main peak are likely a result of profile jitter as an excess on one side of the peak is always accompanied by a deficit on the other. The more subtle long-term changes in the precursor and tail components appear to be genuine state changes in the pulsar emission. The precursor component changes are weakly correlated with variations in the spin-down rate at zero-lag while changes in the tail lag the spin-down rate by $\sim 500$ d.
@@
Identified by Brook et al. (2016) as a variable pulsar, the Gaussian-like profile of PSR J1359--6038 often displays periods of excess emission on the right-hand slope that are correlated with transient decreases in spin-down rate that gradually decay back toward a steady state. An extreme broadening of both the left and right hand sides of the profile coinciding with a sudden $\sim 0.14\%$ drop in spin-down occurred between MJD 56450-56650 (Brook et al., 2016). Despite the much larger overall alteration in pulse shape, the spin-down rate change associated with this event is of comparable amplitude to those that coincide with the less substantial profile shape changes.
@@
This is a $0.3$ s orthogonal rotator that was found to display ‘on/off’ sub-pulse modulation across both its the main and inter pulse (Weltevrede et al., 2007). Like PSR J1057--5226, the spin-down rate of PSR J1705--1906 varies slowly over time, with an apparent, dominant quasi-period of $\sim$2645 d. Our profile variability map does not recover any shape variations that are substantially different from fluctuations in the off-pulse noise, suggesting the rapid state switching every 10.4 rotations identified by Weltevrede et al. (2007) must be stable over long time scales. This is unlike other state switching pulsars, where the profile shape changes recovered in the profile variability maps reflect the pulsar preferentially spending more time in one emission state than the other.
@@
This $1.4$ s pulsar was noted by Johnston et al. (2021) as switching between two distinct emission states: a ‘weak’ state where the profile consists of three approximately equal intensity sub-components and a ‘bright’ mode where the central component is up to an order of magnitude brighter than in the weak-state. Our spin-down model displays small variations with time of up to $\delta\dot{\nu}\sim 0.14\%$, albeit with large uncertainties, that is only weakly correlated with changes in profile shape. A small glitch with an amplitude of $\Delta\nu_{g}/\nu=11.21(5)\times 10^{-9}$ and change in spin-down rate of $\Delta\dot{\nu}_{g}/\dot{\nu}=3.0(2)\times 10^{-3}$ occurred around MJD 58570. Curiously, the profile variability map shows the bright state appears to have been more prevalent in observations taken before MJD 58750. It is unclear if the two phenomena are related since the bright-mode deficit lags the glitch by $\sim 100$ d.
@@
This pulsar was noted by Basu et al. (2024) as displaying alternating fluctuations in intensity between the two peaks of its profile in long-term monitoring observations with MeerKAT. This behaviour appears to be a variant of that identified in the Lyne et al. (2010) sample, where the apparent switching between profile component intensities is due to the pulsar preferentially emitting from one component over the other (see the single-pulse data in Figure 5 of Basu et al. 2024). The switching in component intensities is marginally visible in our profile variability map. Our spin-down timeseries reveals the pulsar initially displayed clear quasi-periodic variability with two peaks in the Lomb-Scargle periodogram at 157 d and 314 d. This behaviour was seemingly interrupted by a glitch that occurred around MJD 58236, after which the quasi-periodicity becomes less defined, which may partially explain the lack of correlation with the emission state changes.
@@
The main peak and trailing shoulder components of this pulsar were noted by Basu et al. (2024) as displaying correlated fluctuations in intensity, where an increase in the emission peak corresponded to a deficit in trailing shoulder emission, and vice-versa. No spin-down variability was reported in their analysis of PSR J1741--3927, however this appears to have been a result of the short timing baseline of the MeerKAT Thousand Pulsar Array, as our $\dot{\nu}$ timeseries shows a clear long-term fluctuation over our 10 yr of Murriyang observations. The profile variability map that we generated also shows the main peak undergoing clear, rapid variations on a timescale of a few hundred days, though we do not see similar anti-correlated changes in the trailing shoulder component. Instead, both the leading and trailing shoulders appear to undergo a secular increase in intensity that is correlated with the long-term change in $\dot{\nu}$ at a 1000 d lag.
@@ 4.2 New pulsars with both profile and spin-down variations
Alongside the known variable pulsars, we have discovered another 28 pulsars that show fluctuations in both profile shape and spin-down rate. Here we summarise the kind of variations detected in the radio emission and spin-down properties of these pulsars.
@@
This pulsar displays a broadly Gaussian shaped radio profile, albeit with many blended sub-components. Its spin-down timeseries in Figure 5 exhibits slow $\delta\dot{\nu}/|\dot{\nu}|\sim 0.72$% fluctuations over 23.5 yr of timing. This is contrasted by the profile variability map, which shows rapid changes across much of the profile peak that appear somewhat uniform across the smaller sub-components. Our cross-correlation analysis returns no substantial positive or negative correlation between the profile variability map and the spin-down timeseries.
@@
Similar to PSR J0742--2822, this pulsar displays a complicated relationship between its highly variable profile shape and spin-down rate. The spin-down rate shows rapid changes with time, with quasi-periods of 456 d and 177 d. The entire profile is affected by shape variations that appear strongest across the peak, which also vary on short timescales. Emission and spin-down processes appear to be uncorrelated with one another similar to the pre-glitch behaviour of PSR J0742--2822.
@@
PSR J1327--6222 is another pulsar that displays rapid, albeit seemingly uncorrelated, variations in both its radio profile shape and spin-down rate over time. The profile shape changes appear in both components of the blended, two-peaked profile with no clear quasi-stable states that the pulsar switches between. Given the rapid changes in both profile shape and spin-down, it is possible our approximately monthly observing cadence is unable to fully resolve any underlying quasi-stable states. These somewhat random profile shape changes are interspersed with brief periods of enhanced emission from a bump in the trailing shoulder, the two most prominent of which occurred around MJD 58915 and MJD 59768.
@@
This pulsar displays slow changes in spin-down rate with an amplitude of $\delta\dot{\nu}/|\dot{\nu}|\sim 0.12$%. However its variable profile shape fluctuates on a much shorter month-long timescales. The observed profile variations appear as increases in width and are most prominent in the trailing edge of the overall Gaussian shape. These changes in the trailing edge are moderately correlated with the spin-down timeseries. Inspection of the profile variability map shows the enhanced emission from this part of the profile appears more prevalent with the pulsar is in a lower spin-down state.
@@
This is a 0.18 s pulsar that has undergone four glitches since the beginning of our timing programme in 2007. The spin-down timeseries of PSR J1453--6413 is largely flat, except for two noteworthy fluctuations in $\dot{\nu}$, one that took place between MJD 57057-58645 and another that began around MJD 59182 and is ongoing. These events appear to be associated with excess emission from the leading bump in the pulse profile, though our correlation analysis suggests there is only a weak positive correlation between the variability map at these pulse phases and the spin-down timeseries. The more rapid profile shape changes that occur in the leading edge of the profile peak are uncorrelated with the changes in $\dot{\nu}$. It is unclear if these profile/spin-down events are related to the small amplitude glitches that took place several hundred days prior to their onset.
@@
PSR J1456--6843 is an example of a pulsar where changes in profile shape are not associated with variations in $\dot{\nu}$. It has a complicated, albeit symmetric pulse profile that appears comprised of multiple blended components. Our variability map shows these components each varying randomly over time, which according to the correlation map is decoupled from the substantially slower $\delta\dot{\nu}/|\dot{\nu}|\sim 0.75$% change in spin-down rate over time.
@@
This pulsar has a triple-component radio profile, with low intensity precursor and trailing components and a bright central component that has a ‘shoulder‘ on its trailing edge. All three components display shape variations different timescales. The precursor and trailing components both show simultaneous excess and suppressed emission that is correlated with the changing spin-down rate of the pulsar. Emission from the leading edge of the central component varies on a similar timescale, but lags changes in the precursor/trailing components by $\sim$400 d. The shoulder on the central component displays more rapid variations, yet appears to preferentially emit in a weak state in concert with the precursor and trailing components. Changes across the central component peak are anti-correlated with the varying spin-down rate and are out of phase with changes in the other components by half a cycle. While the spin-down rate of PSR J1559--4438 visually appears to display both long and short timescale variations, with the more rapid changes appearing quasi-periodic, no substantial peaks are detected in the corresponding Lomb-Scargle periodogram.
@@
PSR J1745--3040 is a bright gamma-ray pulsar with a 0.37 s spin period. It has a two-component radio profile with the brighter trailing component consisting of a bright peak and a leading bump. Both components display substantial epoch-to-epoch changes in intensity in the profile variability map that are uncorrelated with the slow, secular changes in $\dot{\nu}$ over time. This variability appears to be a result of both our short observation times and emission state switching in the single pulses of PSR J1745--3040, behaviour that is also seen in similarly bright pulsars such as PSRs J1048--5832 and J1243--6423. In Figure 11 we present a stack of single pulses detected by Murriyang, where the pulsar clearly switches between on and off emission states. When the pulsar is in the ‘on‘ state, we detect radio emission from both profile components, occasionally interspersed with very bright bursts from the main peak of the second component. Whereas the ‘off‘ state appears to be extended nulls, where no radio emission is detected down to the sensitivity limit of the instrument. A simple comparison between the mean on-pulse flux density and the off-pulse RMS reveals an approximate nulling fraction of 0.51 during the observation presented in Figure 11. Unlike the other pulsars among our sample that display similar behaviour, there does not appear to be a strong association between for the pulsar preferentially emitting in one emission state over the other and the changes in $\dot{\nu}$ over time. This may however, simply be due to the much lower frequency fluctuation timescale of $\dot{\nu}$ and relatively high nulling fraction of this pulsar.
@@
The radio profile of this pulsar consists of a blended, two-component profile that displays rapid fluctuations in intensity confined to the brighter, leading peak. These intensity changes are mildly correlated with the slowly varying spin-down rate. There is a persistent emission deficit that coincided with the pulsar being in a high spin-down state between MJD 58122-59248.
@@
Similar to PSR J0742--2822, this pulsar also displays both rapid variations in profile shape and spin-down rate with a low Q-factor quasi-period of $\sim$157 d. The shape changes affect the entire profile to some degree, but are strongest at the peak. Some of this variability may therefore be the result of profile jitter. However, the shape changes are correlated with the rapidly varying spin-down rate, albeit with a $\sim$50 d lag. No glitches have been detected in the timing of PSR J1752--2806 to date.